use synthetic substitution to evaluate the polynomial for the given value
p(x)=25x²-16 for x=4/5

Answer by tutorcecilia(2144)

p(x)=25x²-16 for x=4/5
=25(4/5)^2-16
=25(16/25)-16
=16-16
=0
.
so,
P(4/5)=0